OpenStudy (anonymous):

4(x - 6) = -2(x - 3)

OpenStudy (anonymous):

4x - 24 = -2x + 6 6x = 30 x= 5

OpenStudy (anonymous):

4(x-6)= -2(x-3) 4x-24=-2x+6 6x=30 x=5